You own $17,390 of Opsware, Inc. stock that has a beta of 3.88. You also own $26,790 of Lowe's companies (beta = 1.72) and $2,820 of New York Times (beta = 0.70). Assume that the market return will be 9 percent and the risk-free rate is 3 percent.

What is the market risk premium?

What is the risk premium of each stock? (3 answers)

What is the risk premium of the portfolio?
Answer #1

Calculation of Risk Premium:

Risk Premium = Market Return rate (Rm)- Risk Free rate (Rf)

As given in the question:

Rm= 9

Rf= 3

Risk Premium = 9-3 = 6%

Calculation Risk Premium of Each Stock:

Risk Premium of each stock= Stock Beta x (Rm-Rf)

Stock Beta Risk Premium= Rm-Rf Stock Risk Premium=
(Stock Beta x Risk Premium)
Opsware, Inc. 3.88 6 23.28
Lowe's companies 1.72 6 10.32
New York Times 0.7 6 4.2

Calculation Risk Premium of Portfolio:

Stock Stock Risk Premium=
(Stock Beta x Risk Premium)
Value Percentage of Portfolio Weighted Average of Stock Risk Premium
Opsware, Inc. 23.28 17390 37.00% 8.6136
Lowe's companies 10.32 26790 57.00% 5.8824
New York Times 4.2 2820 6.00% 0.252
Total 47000 Total 14.7480

Therefore, the Risk premium of Portfolio is 14.7480%